How much of a liquid with a 30% alchohol content and how much of a liquid with a 100% alcohol content should be added to end up with 42 milliiters of with a 50% alcohol content?
Answer by Nate(3500)

[ 100%(x) + 30%(42 - x) ]/[42] = 50%
[ x + 0.3(42 - x) ]/42 = 0.5
[ x + 12.6 - 0.3x ]/42 = 0.5
[ 0.7x + 12.6 ]/42 = 0.5
0.7x + 12.6 = 21
0.7x = 8.4
x = 12
42 - x = 30
12mL of 100% and 30mL of 30%